OpenStudy (anonymous):

Prove that any two medians of a triangle divide each other in the ratio 2:1.

OpenStudy (anonymous):

You know that this is a Mathematical question, wright?

OpenStudy (anonymous):

yaa i knew by mistake instead of posting in geometry i posted it in geography but if u know it u can answer

OpenStudy (anonymous):

1) Let the triangle be ABC; E be the mid point of CA and F be the mid point of AB; Join the medians BE and CF; let them intersect at G; also join EF. 2) Consider triangles, GFE and GBC: i) <FGE = <BGC [Vertically opposite angles] ii) <GFE = <GCB [Alternate interior angles are equla; In a triangle line joining the midpoints of any two sides is parallel to the third side and half of the third side; as such here, EF is parallel to BC and half of BC] iii) Hence triangle GEF is similar to triangle GBC. [AA similarity axiom]. 3) Hence of the above, the corresponding sides are in proportion; that is: GE/GB = GF/GC = EF/BC But EF/BC = 1/2 [Since EF is half of BC from the reason given in 2-ii] ==> GE/GB = GF/GC = 1/2; Thus the centroid G divides each median in the ratio 2:1

OpenStudy (anonymous):

"In a triangle line joining the midpoints of any two sides is parallel to the third side and half of the third side." Is it not the mid point theorem